
In the beginning of the night we could not select the option "object" from the CHIRON software. Manuel Hernandez, Rodrigo Hernandez and Marco Bonati (by phone) help us to solve the problem.
Tremor at 02:39 UT, 5.8 Mw, 35 Km SE of Alto del Carmen. We checked the telescope and everything was ok.
At 05:40 we noted that the headers of the spectra were too short, with a lot of missing information. We checked last days data and noted that this change in the headers started on monday 03rd.
At 08:05 UT we lost the pointing, it was recovered at 08:30 UT.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2018-09-06 | 9.0 | 0.0 | 1.5 | 0.0 | 0.0 | 10.5 |
Links
[1] mailto:thenry@astro.gsu.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7184/delete?destination=print/content/enr-ctio-15-m-2018-09-06
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/7184/edit?destination=print/content/enr-ctio-15-m-2018-09-06
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/14472?destination=print/content/enr-ctio-15-m-2018-09-06
None.